A 2 kg nitrogen gaseous in a piston-cylinder receives heat at a constant pressure of 350 kPaa. The gas temperature increases from 320
1
Expert's answer
2021-07-12T03:36:44-0400

heat at a constant pressure

"Q = mCp \triangle t"


"Q =2 \times 1.23 \times 320 = 787.2 \frac{kJ}{kg}" ..."C_p = 1.23" for "N_2"


Work done :


"W = Q + \triangle U = 787.2+320 = 1107.2 kW"